OpenStudy (anonymous):

I am bit confused about inverse trig functions. Can you manipulate them like this: e.g. arccos (0.6-sin x)=arcsin (cosx-0.2) Is it the same as if I write sin (0.6-sin x)=cos (cosx-0.2)?

OpenStudy (nincompoop):

first, why do you think that you can?

OpenStudy (anonymous):

I don't know, because I thought it could be like other reciprocal functions?

OpenStudy (anonymous):

you ok

OpenStudy (trojanpoem):

You can't simply.However, to get rid of either the arccos or arcsin take cos or sin to both sides: arccos (0.6-sin x)=arcsin (cosx-0.2) cos(arccos(0.6-sinx)) = cos((arcsin(cosx-0.2)) 0.6 - sinx = cos(arcsin(cosx-0.2)) As we know : Cos^2x + Sin^2x = 1 cosx = sqrt(1-sin^2x) 0.6 - sinx = sqrt(1-sin^2(arcsin(cosx-0.2)) 0.6 - sinx = sqrt(1-(cosx-0.2)^2) (0.6-sinx)^2 = (1- (cosx-0.2)^2)

OpenStudy (loser66):

Can you take a snapshot of the original problem? I don't get the question.

OpenStudy (trojanpoem):

@Loser66, He thinks that he can flip the the arcsin with cos, and arccos with sin by taking cos, sin to each site separately without involving the other side.

OpenStudy (anonymous):

@TrojanPoem, thanks!! very clear steps. i understand it now!

OpenStudy (trojanpoem):

You're welcome.